An electron is accelerated from rest through a potential difference is 15V.What is the change in electric potential energy of the electron?


1
Expert's answer
2022-03-10T11:16:07-0500

The change in potential energy is the charge of electron times potential difference:


"\Delta E_p = 1.6\times 10^{-19}C\cdot 15V = 24\times 10^{-19}J"